OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 4) and (5, 1) B. (10, 6) and (1, 5) C. (10, 6) and (5, 1) D. (10, 5) and (5, 1)

OpenStudy (anonymous):

@iambatman

OpenStudy (k_lynn):

Ordered pairs are in (x,y) order. So if you plug in 10 for x in the equation, and 4 for y, then you can see if the equation is correct.

OpenStudy (k_lynn):

So once you plug those in, the equation should look like this. \[4=10-4\]

OpenStudy (k_lynn):

But that's not true so A. isn't the answer. \[4\neq10-4\]

OpenStudy (k_lynn):

You can do that for all of the options until you get one that works. Does that make sense, or should I lead you through it some more?

OpenStudy (anonymous):

yes please because I really dont understand this

OpenStudy (k_lynn):

Ok. So, the ordered pairs are in the order x,y. So x comes first and then comes y. So if you have an ordered pair of (7,14), x=7 and 14=y. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its not D

OpenStudy (anonymous):

@AlexandervonHumboldt2 @AMYCARTER

OpenStudy (anonymous):

Okay let's try B now @verita1959! We're going to substitute 10 for x and 6 for y! \[y=x-4\] ← this becomes this ↓ \[6=10-4\] Now tell me if this is a true statement!

OpenStudy (anonymous):

I dont think it is true but I am not sure I am so confused

OpenStudy (anonymous):

Is \[10-4\] equal to 6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ignore what I just posted lol the equation becomes this ↓ \[5=1-4\] Is this a true statement @verita1959?

OpenStudy (anonymous):

yes@AMYCARTER

OpenStudy (anonymous):

@AMYCARTER

OpenStudy (anonymous):

Okayy so \[1-4\] is equal to 5? I don't really agree @verita1959

OpenStudy (anonymous):

Do we agree that B is not a solution @verita1959? Are you still there?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

so it has to be C?

OpenStudy (anonymous):

It's fine! Now let's move on to C because I have a good feeling about this one! Substitute x,y for 10,6 ↓ We did this earlier remember? \[6=10-4\] We know THAT is true! Now substitute x,y for 5,1 \[1=5-4\] Is C ur answer?

OpenStudy (anonymous):

yes pretty much

OpenStudy (anonymous):

Great job @verita1959! I hope I helped you today :)

OpenStudy (anonymous):

Thank you so much for your help

OpenStudy (anonymous):

np @verita1959 :)

OpenStudy (k_lynn):

Great explanation @AMYCARTER and thanks for taking over, too. I had to go spur of the moment.